OpenStudy (anonymous):

Find all the values of r such that y=Ax^r satisfies the equation: y''+(x^-1)y'-y^3=0 Give a reason why other solutions must exist for this equation.

myininaya (myininaya):

\[y=Arx^{r-1}. y'=Ar(r-1)x^{r-2}, y''=Ar(r-1)(r-2)x^{r-3}\] plug in

OpenStudy (anonymous):

First take the first and second derivatives of your function y with respect to x and plug them into the second order differential equation

OpenStudy (anonymous):

\[r(r-1)Ax^{r-2} + rAx^{r-2}-A^{3} x^{3r} =0\]

OpenStudy (anonymous):

Myininaya the function should be \[y=Ax^r\]

OpenStudy (anonymous):

\[y \prime= rAx ^{r-1}\]

OpenStudy (anonymous):

\[y \prime \prime = r(r-1)A x^{r-2}\]

myininaya (myininaya):

lol

myininaya (myininaya):

thats what i get for going over the speed limit

OpenStudy (anonymous):

hahaha your response was quick

myininaya (myininaya):

my y'' looked to nasty to plug in i felt sorry for him but now i don't

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Here it is with the values plugged in, where do I go from here? \[r(r-1)Ax^{r-2} + rAx^{r-2}-A^{3} x^{3r} =0\]

OpenStudy (anonymous):

Well, A= 0 clearly works...:-)

OpenStudy (anonymous):

We could try adding the coefficient of \[A x^{r-2}\]

OpenStudy (anonymous):

To simplify the first two terms:

OpenStudy (anonymous):

\[A x^{r-2}[r^{2} - r + r]\]

OpenStudy (anonymous):

\[Ar ^{2} x^{r-2}\] - \[A^{3} x^{3r} = 0\]

OpenStudy (anonymous):

Now we have to think lol... Although one of the obvious answers f

OpenStudy (anonymous):

for the other solutions will be that y = 0 is a solution